How much work is done on a table pulled 5m by a force of 30N at an angle of 60° above the horizontal?


1
Expert's answer
2021-04-26T16:48:37-0400

The work done is "A = F \cdot s \cdot \cos \alpha = 30 N \cdot 5 m \cdot \cos 60^{\circ} = 75 J".